A bullet traveling horizontally at a speed of 370m/s hits a board perpendicular to its surface, passes through and emerges on the other side at a speed of 220m/s. If the board is 3.90 thick, how long does the bullet take to pass through it?
1
Expert's answer
2013-02-11T10:28:33-0500
The average speed of the bullet while it penetrates the board is

V = (370[m/s] + 220[m/s])/2 = 295[m/s].

Assuming that the velocity of the bullet decreases linearly, it takes

T = 3.90[cm]/295[m/s] = 0.039[m]/295[m/s] = 1.322·10^(-4)[s]

seconds the bullet to pass through the board.